OpenStudy (bayleejadee):

The equation A = p + prt is used to find the accumulated amount, A, of an investment with principal, p, a certain interest rate, r, and over time, t. What is the value for t, after the equation is rearranged?

OpenStudy (bayleejadee):

t=A-p over pr t= A+p over pr t= A-1 over r t = A − p − pr

OpenStudy (bayleejadee):

@royalranger @Hero @shirleyarnett

OpenStudy (royalranger):

You have any idea where to start?

OpenStudy (bayleejadee):

Honestly? Not a clue

OpenStudy (royalranger):

Alright so we want to solve for t, therefore we must isolate it. a = p + prt Lets begin by subtracting p: a-p = prt What do we do next in order to get t by itself?

OpenStudy (royalranger):

Example: 

OpenStudy (bayleejadee):

Divide by r? I'm kind of just winging it here

OpenStudy (bayleejadee):

Sorry I took so long to answer by the way, I was working on another problem.

OpenStudy (royalranger):

Yes you could divide by r, but if you did, you would be left with: \[\frac{ a-p }{ r } = pt\] then you could have to divide by p to get t by itself. So instead, why don't you try dividing by pr to get t by itself.

OpenStudy (bayleejadee):

So, It would be the first option then?

OpenStudy (royalranger):

Yep

OpenStudy (bayleejadee):

Thank you!

OpenStudy (royalranger):

Yep